OpenStudy (anonymous):

what is the second derivative of the parametic function x = sqrt(t), y =8t-1, t= 4, if the first derivative is equal to 32? ( I did the math for the first deriv., but I cannot get the second one.)

OpenStudy (anonymous):

Hello!

OpenStudy (anonymous):

did you get first derivative: \[=16\sqrt{t}\] ?

OpenStudy (anonymous):

No, I got the derivative as 8/(1/2t^1/2)

OpenStudy (anonymous):

or 8/(1/4)

OpenStudy (anonymous):

I got for second derivative: \[d/dt (dy/dx)/(dx/dt)\] \[d(16\sqrt{t})dt/(dx/dt)=16\]

OpenStudy (anonymous):

But i don't think I did it right, even if 32 is the right answer

OpenStudy (anonymous):

i just went with definition of second derivative and fact that y=8t-1=8x^2-1 (put the value of x=sqrt t)

OpenStudy (anonymous):

I had 1/2t^(-1/2) as the dx/dt for the first derivative

OpenStudy (anonymous):

I got 16.... so I missed *2 somewhere... let me check

OpenStudy (anonymous):

and y = 8 for dy/dt

OpenStudy (anonymous):

\[dy/dx=(dy/dt)/*dx/dt)=8/(1/2\sqrt{t})=16\sqrt{t}\] right?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

yes, but I

OpenStudy (anonymous):

i missed typed denominator... shouldn't be any "*"

OpenStudy (anonymous):

wouldn't it be 16/ sqrt(t)?

OpenStudy (anonymous):

\[(\sqrt{t})'=1/(2\sqrt{t})\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

But even with 16/sqrt(t), you would but in 4 and get 16/2 right? Which is 8 though

OpenStudy (anonymous):

Hey visitor! Sign up for Open Study, it's awesome!!

OpenStudy (anonymous):

let's try one more time: \[dy/dx=(dy/dt) / (dx/dt)=8/(1/2\sqrt{t})=16\sqrt{t}\]

OpenStudy (anonymous):

you mean 16/sqrt(t)

OpenStudy (anonymous):

no. there is no division.

OpenStudy (anonymous):

????

OpenStudy (anonymous):

okay, so the 1/2 multiplied to sqrt(t) doesn't get moved to the top

OpenStudy (anonymous):

\[8/(1/(2*\sqrt{t}))=8*2*\sqrt{t}\]

OpenStudy (anonymous):

so, if 16sqrt(t), then 16* sqrt(4) = 16*2 is 32. cool. we got it. Now, the second derivative though?

OpenStudy (anonymous):

second derivative\[=d/dt (dy/dx) : dx/dt\] \[y=8t-1=8x ^{2}-1\] (use substitution from first eq) agreed?

OpenStudy (anonymous):

where did the x come from?

OpenStudy (anonymous):

\[x=\sqrt{t}\] given

OpenStudy (anonymous):

okay, yep agreed

OpenStudy (anonymous):

so, 16x?

OpenStudy (anonymous):

yes it is!!!! thanks for all your help!:D

OpenStudy (anonymous):

awesome!